OpenStudy (andijo76):

In a sample of 7 cars each car was tested for nitrogen-oxide emissions (in grams per mile) and the following results are obtained 0.18, 0.16, 0.08, 0.06,0.17,0.12,0.15 assuming that this sample is representative of the cars in use. cosntruct a 98% confidence interval estimate of the mean amount of nitrogen=oxide emissions for all cars. if the epa requires that nitrogen oxide emissions can be less than 0.165g.mi we can safely conclude that requirement is being met A)what is the confidence interval estimate of the mean amount of nitrogen-oxide emissions for all cars ___g/mi

OpenStudy (anonymous):

I have this same question! lol

OpenStudy (kropot72):

The first step is to calculate the sample mean by adding the seven values and dividing by 7: \[\bar{x}=\frac{0.18+0.16+0.08+0.06+0.17+0.12+0.15}{7}=\frac{0.92}{7}=0.13143\]

OpenStudy (anonymous):

@kropot72, then how do you get the __g/mi <u< __g/mi ??

OpenStudy (kropot72):

The next step is to calculate the sample standard deviation, s.

OpenStudy (anonymous):

not sure how to do that..

OpenStudy (kropot72):

You can use a calculator with statistical functions to calculate the sample standard deviation, s.

OpenStudy (andijo76):

x= i got a totally different answer than you did

OpenStudy (andijo76):

and the sd i think is 0.046

OpenStudy (anonymous):

how did you get the sd @andijo76

OpenStudy (kropot72):

@andijo76 what did you get for the sum of the seven data values?

OpenStudy (andijo76):

i found a website back when i was doing pre algebra that helps me figuring out some calculations

OpenStudy (andijo76):

the sum i got was 0.842/7=0.1203

OpenStudy (anonymous):

does anyone know how to input this into the Ti84?

OpenStudy (kropot72):

@andijo76 The correct sum of the seven data values is 0.92.

OpenStudy (andijo76):

im trying to figure out how i got that lol

OpenStudy (andijo76):

is the sd right tho

OpenStudy (kropot72):

My calculator gives the value for the sample standard deviation of s = 0.4634

OpenStudy (anonymous):

@kropot72 how did you input the sd in the calc?

OpenStudy (kropot72):

My calculator is a Casio fx-82 AU PLUS. I just followed the instructions in the User's Guide.

OpenStudy (andijo76):

im sorry you are correct i dont know what i did but i fixed it i think i rounded one decimal or something

OpenStudy (andijo76):

there is a website that gives you sd i will paste what it said

OpenStudy (andijo76):

Simplify the result. σ=0.046344719225 The standard deviation should be rounded to one more decimal place than the original data. If the original data were mixed, round to one decimal place more than the least precise. 0.046

OpenStudy (anonymous):

@andijo76 can you paste the link to the website?

OpenStudy (andijo76):

www.mathway.com it costs 20 a month though i keep it for my kids it helps you figure out things step by step

OpenStudy (kropot72):

The final calculation uses the following formula: \[\bar{x}-2.576\frac{s}{\sqrt{n}}<\mu<\bar{x}+2.576\frac{s}{\sqrt{n}}\] Substituting values gives the confidence interval for the population mean as follows: \[0.13143-2.576\frac{0.046}{\sqrt{7}}<\mu<0.13143+2.576\frac{0.046}{\sqrt{7}}\]

OpenStudy (anonymous):

where did you get the 2.576?

OpenStudy (kropot72):

The value 2.576 is determined by an inverse normal technique. However it can be found in tables for use in calculating confidence intervals.

OpenStudy (anonymous):

ok

OpenStudy (andijo76):

i know i am going to sound stupid but how do i figure _g/mi<u<_g/mi was that the answer

OpenStudy (kropot72):

When you have calculated the two values for the confidence interval, just put the lower value where the first underscore is and the higher value where the right hand underscore is.

OpenStudy (andijo76):

i got for the first one -44.6558 so i am thinking it is i just have to put round and put it in 3 decimal places

OpenStudy (anonymous):

so is this the answer then: .0866 g/mi< .1762 g/mi ?! @kropot72

OpenStudy (kropot72):

Looks correct to me :)

OpenStudy (andijo76):

yeah i forgot the point haha

OpenStudy (andijo76):

it amazing how that works hey

OpenStudy (kropot72):

It is a very powerful technique!

OpenStudy (andijo76):

thank you so much for your help

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

lol great! thanks!!